# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v1

@@ 3 Differentiating a Parametric Fixed Point
where $A_{\lambda}\colon\mathbb{R}^{d}\rightrightarrows\mathbb{R}^{d}$ and $B_{\lambda}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ are multi-valued and single-valued maximal monotone operators respectively. These types of problems are at the core of convex analysis and can cover a number of optimizations problems including minimization problems as well as variational inequalities and saddle points problems. It is a standard fact that (6) can be rewritten as the equation
@@
where $f_{\lambda}=\mathbb{E}\hat{f}_{\lambda}(\cdot,\xi)$ is $L$-smooth and convex while $g_{\lambda}$ is convex lower semicontinuous extended-real valued. This can be cast into (2) by setting $\eta\in\left[0,2/L\right[$, $\hat{T}(w,\lambda,\xi)=w-\eta\hat{\nabla}f_{\lambda}(w,\xi)$ and $G(w,\lambda)=\operatorname{Prox}_{\eta g_{\lambda}}(w)$ with $\operatorname{Prox}_{h}(x)=\operatorname*{arg\,min}_{y}(h(x)+(1/2)\lVert x-y\rVert^{2})$ being the proximal operator of $h$. Several machine learning problems can be written in form (7), e.g. LASSO, elastic net, (dual) SVM, where $g_{\lambda}$ is not smooth.
@@
The first relation above shows that $w^{\prime}(\lambda)\in\mathbb{R}^{d\times p}$ is a fixed point of the linear contraction $X\mapsto\partial_{1}\Phi(w(\lambda),\lambda)X+\partial_{2}\Phi(w(\lambda),\lambda)$.
Here, dealing with the nonsmooth case, we will mimic the above formulas. The following will be a crucial assumption of our analysis.
@@
Let $O_{\Lambda}\subset\mathbb{R}^{m}$ be open, with $\Lambda\subset O_{\Lambda}$ be nonempty closed and convex.
@@
    $\Phi\colon\mathbb{R}^{d}\times O_{\Lambda}\to\mathbb{R}^{d}$ is definable and a continuous selection of the $L$-Lipschitz smooth definable mappings $\Phi_{1},\dots,\Phi_{m}$ and we set $D_{\Phi}\colon\mathbb{R}^{d}\times O_{\Lambda}\rightrightarrows\mathbb{R}^{d\times(d+p)}$,
    $D_{\Phi}(u,\lambda)=D_{\Phi,1}^{s}(u,\lambda)\times D_{\Phi,2}^{s}(u,\lambda),$ (9)
    where $D_{\Phi,k}^{s}(u,\lambda)=\operatorname{co}(\{\partial_{k}\Phi_{i}(u,\lambda)\,|\,i\in I_{\Phi}(u,\lambda)\})$, with $k=1,2$, are the partial conservative derivatives.
@@
    For every $(u,\lambda)\in\mathbb{R}^{p}\times O_{\Lambda}$, $\lVert D_{\Phi,1}(u,\lambda)\rVert_{\sup}\leq q<1$.
The results in the previous section and Lemma 3 and Lemma 4 in (Bolte & Pauwels, 2021), ensure that $D_{\Phi}$, as defined in (9), is a conservative derivative of $\Phi$. Moreover, recalling (3), it is easy to see that item (ii) in the above assumption ensures that $\Phi(\cdot,\lambda)$ is a $q$-contraction and hence that there exists a unique fixed point of $\Phi(\cdot,\lambda)$ that we will denote by $w(\lambda)$. Finally, if $A\in D_{\Phi,1}(u,\lambda)$, we have $\lVert A\rVert<1$ and hence $I-A$ is invertible. Thus, mimicking what happens for the smooth case (see (8)) one defines
$D_{w}^{\operatorname{imp}}(\lambda)=(I-D_{\Phi,1}(w(\lambda),\lambda))^{-1}D_{\Phi,2}(w(\lambda),\lambda)$ (10)
@@
where $\mathrm{fix}[D_{\Phi}(u,\lambda)]$ is the unique fixed “point” of the map $\mathcal{X}\mapsto D_{\Phi,1}(u,\lambda)\mathcal{X}+D_{\Phi,2}(u,\lambda)$, acting between compact sets of $d\times p$ matrices. In (Bolte et al., 2022) it is proved that if $\Phi$ is path differentiable and Assumption 3.1(ii) holds, the set-valued mappings $D_{w}^{\operatorname{imp}}$ and $D_{w}^{\operatorname{fix}}$ are both conservative derivatives of $w(\lambda)$ and $D_{w}^{\text{imp}}(\lambda)\subset D_{w}^{\operatorname{fix}}(\lambda)$. We note that the product structure (9) of $D_{\Phi}$ allows us to express $D_{w}^{\operatorname{imp}}$ and $D_{w}^{\operatorname{fix}}$, originally introduced in (Bolte et al., 2022), in terms of Minkowski set operations. This will facilitate deriving convergence rates.
3.1 yields the following lemma through an application of Lemma 2.5 (the proof is in Appendix B).
@@
Under 3.1(i), for every $\lambda\in\Lambda$, there exist $R_{\lambda}>0$ such that for $k=1,2$
$\forall\,u\in\mathbb{R}^{d}\colon\ \operatorname{\it{e}}(D_{\Phi,k}(u,\lambda),D_{\Phi,k}(w(\lambda),\lambda))\leq C_{\lambda}(u)\lVert u-w(\lambda)\rVert,$
@@
[…]
and $M_{\lambda}\!:=\max\limits_{i\in[m]}\min\limits_{j\in I(w(\lambda),\lambda)}\lVert\Phi^{\prime}_{i}(w(\lambda),\lambda)-\Phi^{\prime}_{j}(w(\lambda),\lambda)\rVert$.
@@ 4 Deterministic Iterative and Approximate Implicit Differentiation
This method approximates $D_{w}^{\operatorname{fix}}(\lambda)$ through the iterative procedure
$\begin{array}[]{l}\text{for}\;t=1,2\ldots\\ \left\lfloor\begin{array}[]{l}w_{t}(\lambda)=\Phi(w_{t-1}(\lambda),\lambda)\\[3.44444pt] D_{w_{t}}(\lambda)=D_{\Phi,1}(w_{t-1}(\lambda),\lambda)D_{w_{t-1}}(\lambda)+D_{\Phi,2}(w_{t-1}(\lambda),\lambda).\end{array}\right.\\ \end{array}$ (13)
Note that the iteration for $D_{w_{t}}(\lambda)$ is based on the chain rule and results in a conservative derivative of $w_{t}(\lambda)$. Under the product structure of $D_{\Phi}$ in (9), this is the same set-valued iteration studied in (Bolte et al., 2022). We note that if $\Phi(\cdot,\lambda)$ is a $q$-contraction, it holds $\lVert w_{t}(\lambda)-w(\lambda)\rVert=O(q^{t})$.
@@
$\begin{array}[]{l}\text{for}\;k=1,2\ldots\\ \left\lfloor\begin{array}[]{l}D_{w_{t}}^{k}(\lambda)=D_{\Phi,1}(w_{t}(\lambda),\lambda)D_{w_{t}}^{k-1}(\lambda)+D_{\Phi,2}(w_{t}(\lambda),\lambda).\end{array}\right.\\ \end{array}$ (14)
@@
Let $x\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{d}$. The ITD method exploits automatic differentiation to efficiently compute an element of the conservative Jacobian-vector products $D_{w_{t}}(\lambda)^{\top}y$ (in reverse mode) and $D_{w_{t}}(\lambda)x$ (in forward mode). Similarly AID efficiently computes an element in $D_{w_{t}}^{k}(\lambda)^{\top}y$. Thanks to Automatic Differentiation, if $k=t$ the standard implementation of both AID-FP and ITD has a cost in time of the same order of that of computing $w_{t}(\lambda)$. However, while AID-FP only uses $w_{t}(\lambda)$, ITD has a larger $\Theta(t)$ memory cost , since it needs to store the entire optimization trajectory $(w_{i}(\lambda))_{0\leq i\leq t}$.
@@
Let 3.1 hold. For every $\lambda\in\Lambda$, let $R_{\lambda}$ and $M_{\lambda}$ be the quantities defined in Lemma 3.2 and $B_{\lambda}:=\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}$. For every $t,k\in\mathbb{N}$, let $\Delta_{t}=\lVert w_{t}(\lambda)-w(\lambda)\rVert$, $E_{\lambda}(t):=\mathds{1}\{\Delta_{t}>R_{\lambda}\}$ and $\bar{E}_{\lambda}(t)=t^{-1}\sum_{i=0}^{t-1}E_{\lambda}(i)$. Then the following hold.
@@
    If $D_{w_{0}}(\lambda)=\{0\}$, the ITD iteration in (13) satisfies
    $\operatorname{\it{e}}(D_{w_{t}}(\lambda),D_{w}^{\operatorname{fix}}(\lambda))\leq\frac{B_{\lambda}}{1-q}\,q^{t}+\frac{B_{\lambda}+1}{1-q}\Big{(}L+\frac{M_{\lambda}}{R_{\lambda}}\bar{E}_{\lambda}(t)\Big{)}\Delta_{0}\,t\,q^{t-1}.$ (15)
@@
    $\operatorname{\it{e}}(D_{w_{t}}^{k}(\lambda),D_{w}^{\operatorname{fix}}(\lambda))\leq\frac{B_{\lambda}}{1-q}\,q^{k}+\frac{B_{\lambda}+1}{1-q}\Big{(}L+\frac{M_{\lambda}}{R_{\lambda}}E_{\lambda}(t)\Big{)}\frac{1-q^{k}}{1-q}\Delta_{t}.$ (16)
Moreover, if $w_{t}(\lambda)=\Phi(w_{t-1}(\lambda),\lambda)$, then $\Delta_{t}\leq q\Delta_{t-1}\leq q^{t}\Delta_{0}$ and there exists $\tau_{\lambda}\in\mathbb{N}$ such that $E(t)=\mathds{1}\{t<\tau_{\lambda}\}$ and thus $E_{\lambda}(t)\leq\bar{E}_{\lambda}(t)\leq 1$.
To compare the two rates in Theorem 4.1, let $t=k$ and $w_{t}(\lambda)=\Phi(w_{t-1}(\lambda),\lambda)$, so that both AID-FP and ITD have time complexity of the order of computing $w_{t}(\lambda)$. In that situation, since $1-q^{k}=1-q^{t}<q^{-1}(1-q)t$ and $E_{\lambda}(t)\leq\bar{E}_{\lambda}(t)$, the upper bound of AID-FP is always lower than that of ITD. Moreover, if we let $\kappa=(1-q)^{-1}$ to play a similar role to the condition number, we observe that both methods converge linearly: AID-FP as $O(\kappa^{2}e^{-t/\kappa})$, while ITD slightly slower as $O(\kappa te^{-t/\kappa})$. When $t\geq\tau_{\lambda}$, $E_{\lambda}(t)=0$ while $\bar{E}_{\lambda}(t)=\tau_{\lambda}/t$, which might cause a wide difference between the two bounds if $M_{\lambda}/R_{\lambda}$ is large, and such ratio can get arbitrarily large the closer $(w(\lambda),\lambda)$ is to regions where $\Phi$ is not differentiable. Finally, if we replace Lemma 3.2 with the $L$-smoothness of $\Phi$, we essentially recover the same bounds reported by Grazzi et al. (2020), where the terms $E_{\lambda},\bar{E}_{\lambda}$ do not appear.
Note that since the iterations (13) and (14) are set-valued, our analysis works even if the element of the conservative derivative selected at each iteration of AID-FP and ITD changes from one iteration to the other. However, in practice this is rarely the case, instead the same element $\Phi^{\prime}\in D_{\Phi}$ is used at every iteration. This choice yields that both AID-FP and ITD are guaranteed to converge, up to a subsequence, to an element of $D_{w}^{\operatorname{fix}}(\lambda)$.

### v2

@@ 3 Differentiating a Parametric Fixed Point
where $A_{\lambda}\colon\mathbb{R}^{d}\rightrightarrows\mathbb{R}^{d}$ and $B_{\lambda}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ are multi-valued and single-valued maximal monotone operators respectively. These types of problems are at the core of convex analysis and can cover a number of optimizations problems including minimization problems as well as variational inequalities and saddle points problems. It is a standard fact (see (Bauschke & Combettes, 2017)) that (7) can be rewritten as the equation
@@
where $f_{\lambda}=\mathbb{E}\hat{f}_{\lambda}(\cdot,\xi)$ is convex $L$-smooth, while $g_{\lambda}$ is convex lower semicontinuous extended-real valued. This can be cast into (2) by setting $\eta\in\left[0,2/L\right[$, $\hat{T}_{\xi}(w,\lambda)=w-\eta\hat{\nabla}f_{\lambda}(w,\xi)$ and $G(w,\lambda)=\operatorname{Prox}_{\eta g_{\lambda}}(w)$ with $\operatorname{Prox}_{h}(x)=\operatorname*{arg\,min}_{y}(h(x)+(1/2)\lVert x-y\rVert^{2})$ being the proximity operator of $h$. Several machine learning problems can be written in form (8), e.g., LASSO, elastic net, (dual) SVM, where $g_{\lambda}$ is not smooth.
@@
The first relation above shows that $w^{\prime}(\lambda)\in\mathbb{R}^{d\times p}$ is a fixed point of the map $X\mapsto\partial_{1}\Phi(w(\lambda),\lambda)X+\partial_{2}\Phi(w(\lambda),\lambda)$. Here, dealing with the nonsmooth case, we will mimic the above formulas. The crucial assumption of our analysis is the following.
@@
Let $O_{\Lambda}\subset\mathbb{R}^{m}$ be an open set and $\Lambda\subset O_{\Lambda}$ be a nonempty closed and convex set.
@@
    $\Phi\colon\mathbb{R}^{d}\times O_{\Lambda}\to\mathbb{R}^{d}$ is definable and a continuous selection of the $L$-Lipschitz smooth definable mappings $\Phi_{1},\dots,\Phi_{r}$ and we set $D_{\Phi}\colon\mathbb{R}^{d}\times O_{\Lambda}\rightrightarrows\mathbb{R}^{d\times(d+m)}$,
    $D_{\Phi}(u,\lambda){=}D_{\Phi}^{s}(u,\lambda){=}\operatorname{co}(\{\Phi^{\prime}_{i}(u,\lambda)\,|\,i\in I_{\Phi}(u,\lambda)\}).$ (10)
@@
    For all $(u,\lambda)\in\mathbb{R}^{p}\times O_{\Lambda}$, $\lVert D_{\Phi,1}(u,\lambda)\rVert_{\sup}\leq q<1$.
Theorem 2.4 ensures that $D_{\Phi}$, as defined in (10), is a conservative derivative of $\Phi$. Moreover, recalling (4), it is easy to see that 3.1(ii) ensures that $\Phi(\cdot,\lambda)$ is a $q$-contraction and hence that there exists a unique fixed point of $\Phi(\cdot,\lambda)$ that we will denote by $w(\lambda)$. Finally, if $A\in D_{\Phi,1}(u,\lambda)$, we have $\lVert A\rVert<1$ and hence $I-A$ is invertible. Thus, mimicking what happens for the smooth case in (9) one defines
$D_{w}^{\operatorname{imp}}(\lambda){=}\big{\{}(I-A_{1})^{-1}A_{2}\,|\,[A_{1},A_{2}]\in D_{\Phi}(w(\lambda),\lambda)\big{\}}$ (11)
@@
where $\mathrm{fix}[D_{\Phi}(u,\lambda)]$ is the unique fixed “point” of the map $\mathcal{X}\mapsto\mathcal{A}(\mathcal{X})$, where $\mathcal{A}=D_{\Phi}(u,\lambda)$ (see equation (3)), which acts between compact sets of $d\times m$ matrices. In (Bolte et al., 2021) it is proved that if $\Phi$ is path differentiable and Assumption 3.1(ii) holds, the set-valued mappings $D_{w}^{\operatorname{imp}}$ and $D_{w}^{\operatorname{fix}}$ are both conservative derivatives of $w(\lambda)$ and $D_{w}^{\text{imp}}(\lambda)\subset D_{w}^{\operatorname{fix}}(\lambda)$.
3.1 yields the following lemma through a direct application of Lemma 2.5.
@@
Under 3.1(i), for every $\lambda\in\Lambda$, there exist $R_{\lambda}>0$ such that for every $u\in\mathbb{R}^{d}$
$\operatorname{\it{e}}(D_{\Phi}(u,\lambda),D_{\Phi}(w(\lambda),\lambda))\leq C_{\lambda}(u)\lVert u-w(\lambda)\rVert,$
@@
$C_{\lambda}(u):=\begin{cases}L&\text{if }\lVert u-w(\lambda)\rVert\leq R_{\lambda}\\ L+M_{\lambda}/R_{\lambda}&\text{otherwise}\end{cases}$ (13)
and $M_{\lambda}\!:=\max\limits_{i\in[r]}\min\limits_{j\in I_{\Phi}(w(\lambda),\lambda)}\lVert\Phi^{\prime}_{i}(w(\lambda),\lambda)-\Phi^{\prime}_{j}(w(\lambda),\lambda)\rVert$.
@@
Remark 3.3.
Our theoretical analysis requires only that $\Phi$ is definable piecewise smooth and that the inequality in Lemma 3.2 holds for some conservative derivatives of $\Phi$, even if it is not computed according to (10). One such situation occurs for instance when $\Phi$ has the structure of a finite sum, that is, $\Phi=\sum_{i=1}^{n}\Phi^{(i)}$, where each $\Phi^{(i)}$ satisfies Assumption 3.1(i) with corresponding conservative derivative $D_{\Phi^{(i)}}^{s}$. Then, it is clear that $\Phi$ is still definable and piecewise Lipschitz smooth. Moreover, using the properties of conservative derivatives (see Corollary 4 in (Bolte & Pauwels, 2020)), $D_{\Phi}=\sum_{i=1}^{n}D_{\Phi^{(i)}}^{s}$ is a conservative derivatives of $\Phi$. Thus, using the property of the excess (see Lemma B.1(ii)) it directly follows that the inequality in Lemma 3.2, and hence our theory, still holds for such $\Phi$.
@@ 4 Deterministic Iterative and Approximate Implicit Differentiation
This method approximates $D_{w}^{\operatorname{fix}}(\lambda)$ through the following iterative procedure, starting from $w_{0}(\lambda)\in\mathbb{R}^{d}$, $D_{w_{0}}(\lambda)=\{0\}$,
$\begin{array}[]{l}\text{for}\;t=1,2\ldots\\ \left\lfloor\begin{array}[]{l}w_{t}(\lambda)=\Phi(w_{t-1}(\lambda),\lambda)\\[3.44444pt] D_{w_{t}}(\lambda)=D_{\Phi}(w_{t-1}(\lambda),\lambda)\,\begin{bmatrix}D_{w_{t-1}}(\lambda)\\ I_{m}\end{bmatrix},\end{array}\right.\end{array}$ (14)
where we used the definition in (3). Note that the iteration for $D_{w_{t}}(\lambda)$ is based on the chain rule and results in a conservative derivative of $w_{t}(\lambda)$. This is the same set-valued iteration studied in (Bolte et al., 2022). We note that if $\Phi(\cdot,\lambda)$ is a $q$-contraction, it holds $\lVert w_{t}(\lambda)-w(\lambda)\rVert=O(q^{t})$.
@@
$\begin{array}[]{l}\text{for}\;k=1,2\ldots\\ \left\lfloor\begin{array}[]{l}D_{w_{t}}^{k}(\lambda)=D_{\Phi}(w_{t}(\lambda),\lambda)\,\begin{bmatrix}D_{w_{t}}^{k-1}(\lambda)\\ I_{m}\end{bmatrix}.\end{array}\right.\end{array}$ (15)
@@
In practice we do not compute the full set-valued iterations in (14) and (15), but rather we select just one element at each iteration. Moreover, if we let $x\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{d}$, the ITD method can exploit automatic differentiation to efficiently compute an element of the conservative Jacobian-vector products $D_{w_{t}}(\lambda)^{\top}y$ (in reverse mode) and $D_{w_{t}}(\lambda)x$ (in forward mode). Similarly AID can efficiently compute an element in $D_{w_{t}}^{k}(\lambda)^{\top}y$. Thanks to Automatic Differentiation, if $k=t$ the standard implementation of both AID-FP and ITD has a cost in time of the same order of that of computing $w_{t}(\lambda)$. However, while AID-FP only uses $w_{t}(\lambda)$, ITD has a larger $\Theta(t)$ memory cost, since it needs to store the entire optimization trajectory $(w_{i}(\lambda))_{0\leq i\leq t}$.
@@
Let 3.1 hold. For every $\lambda\in\Lambda$, let $R_{\lambda}$ and $M_{\lambda}$ be the quantities defined in Lemma 3.2 and $B_{\lambda}:=\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}$. For every $t,k\in\mathbb{N}$, let $\Delta_{t}=\lVert w_{t}(\lambda)-w(\lambda)\rVert$, $\delta_{\lambda}(t):=\mathds{1}\{\Delta_{t}>R_{\lambda}\}$ and $\bar{\delta}_{\lambda}(t)=t^{-1}\sum_{i=0}^{t-1}\delta_{\lambda}(i)$. Then the following hold.
@@
    The ITD iteration in (14) satisfies
    $\operatorname{\it{e}}(D_{w_{t}}$ $(\lambda),D_{w}^{\operatorname{fix}}(\lambda))\leq\frac{B_{\lambda}}{1-q}\,q^{t}+\frac{B_{\lambda}+1}{1-q}\Big{(}L+\frac{M_{\lambda}}{R_{\lambda}}\bar{\delta}_{\lambda}(t)\Big{)}\Delta_{0}\,t\,q^{t-1}.$ (16)
@@
    $\operatorname{\it{e}}(D_{w_{t}}^{k}$ $(\lambda),D_{w}^{\operatorname{fix}}(\lambda))\leq\frac{B_{\lambda}}{1-q}\,q^{k}+\frac{B_{\lambda}+1}{1-q}\Big{(}L+\frac{M_{\lambda}}{R_{\lambda}}\delta_{\lambda}(t)\Big{)}\frac{1-q^{k}}{1-q}\Delta_{t}.$ (17)
Moreover, if $w_{t}(\lambda)=\Phi(w_{t-1}(\lambda),\lambda)$, then $\Delta_{t}\leq q\Delta_{t-1}\leq q^{t}\Delta_{0}$ and there exists $\tau_{\lambda}\in\mathbb{N}$ such that $\delta_{\lambda}(t)=\mathds{1}\{t<\tau_{\lambda}\}$ and thus $\delta_{\lambda}(t)\leq\bar{\delta}_{\lambda}(t)\leq 1$.
To compare the two rates in Theorem 4.1, let $t=k$ and $w_{t}(\lambda)=\Phi(w_{t-1}(\lambda),\lambda)$, so that both AID-FP and ITD have time complexity of the order of computing $w_{t}(\lambda)$. In that situation, since $1-q^{k}=1-q^{t}<q^{-1}(1-q)t$ and $\delta_{\lambda}(t)\leq\bar{\delta}_{\lambda}(t)$, the upper bound of AID-FP is always lower than that of ITD. Moreover, if we let $\kappa=(1-q)^{-1}$ to play a similar role to the condition number, we observe that both methods converge linearly: AID-FP as $O(\kappa^{2}e^{-t/\kappa})$, while ITD slightly slower as $O(\kappa te^{-t/\kappa})$. When $t\geq\tau_{\lambda}$, $\delta_{\lambda}(t)=0$ while $\bar{\delta}_{\lambda}(t)=\tau_{\lambda}/t$, which might cause a wide difference between the two bounds if $M_{\lambda}/R_{\lambda}$ is large, and such ratio can get arbitrarily large the closer $(w(\lambda),\lambda)$ is to regions where $\Phi$ is not differentiable. Finally, if we replace Lemma 3.2 with the $L$-smoothness of $\Phi$, we essentially recover the same bounds reported by Grazzi et al. (2020), where the terms $\delta_{\lambda},\bar{\delta}_{\lambda}$ do not appear.
The work by (Bolte et al., 2022) also reports a rate for nonsmooth ITD of $O((\sqrt{q}+\epsilon)^{t})$ for arbitrary $\epsilon>0$. However, this rate does not match the best available rate for smooth ITD (Grazzi et al., 2020). Theorem 4.1 (in (16)) fills this gap since it achieves [^3] For any $\epsilon\in[0,1-q]$, $\exists C>0$ such that $tq^{t-1}\leq C(q+\epsilon)^{t}$. an improved rate of $O((q+\epsilon)^{t})$. Moreover, our rate is more explicit, since it does not involve any arbitrary $\epsilon$.
We conclude the section by noting that Theorem 4.1 ensures that the sequence constructed by selecting one element at each iteration in (14) and (15), is guaranteed to converge, up to a subsequence, to the set $D_{w}^{\operatorname{fix}}(\lambda)$.